Two cyclists racing on parrell roads maintain constant speeds of 30 mph and 25 mph. The faster cyclist crosses the finish line one hour before the slower cyclist. How long was the race in miles?
Answer by checkley77(12844)

D=RT
30(T-1)=25T
30T-30=25T
30T-25T=30
5T=30
T=30/5
T=6 HOURS IS THE RACE TIME FOR THE SLOWER CYCLIST.
25*6=150 MILES IS THE RACE DISTANCE.